The difference of the present ages of two brothers is 4 years,5years ago,if the product of their ages was 96,find their ages.
Answer by mananth(16946)

let elder brother's age be x
the youger brother's age be y
..
x-y=4
x=y+4
..
five years ago they were x-5
& y-5
..
(x-5)(y-5)=96
xy-5x-5y+25=96
xy-5x-5y=96-25
xy-5x-5y=71
plug the value of x
(y+4)y-5(y+4)-5y=71
y^2+4y -5y-20-5y=71
y^2-6y-91=0
y^2-13y+7y-91=0
y(y-13)+7(y-13)=0
(y-13)(y+7)=0
y =13 ignore negative value.
x=y+4
x=13+14=17
The ages are 13 , 17 years
CHECK
(17-5)(13-5)=96